OpenStudy (anonymous):

how do you find the angles of an isosolies triangle if all you have is the measures of all of the sides?

OpenStudy (amistre64):

whats the side measures?

OpenStudy (amistre64):

SSS i believe is a valid form

OpenStudy (amistre64):

which one is an isoslese again?

OpenStudy (anonymous):

Let b and s be the length of the base side and the length of either of the two equal sides respectively. Let theta be the angle opposite one of the equal sides. Apply the Law of Sines.\[\frac{s}{\text{Sin}[\theta ]}=\frac{b}{\text{Sin}[\pi -2\theta ]}\] and solve for theta.\[\theta =\text{ArcSec}\left[\frac{2 s}{b}\right] \]